OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. x + y = −3 y = 2x + 2 A. 5/3, 4/3 B. -5/3, -4/3 C. -3/5, -3/4 D. 3/4, 3/5

OpenStudy (anonymous):

i think c

OpenStudy (er.mohd.amir):

ans B x + y =-3 put value of y x +2x+2=-3 3x=-3-2 3x=-5 x=-5/3 put value of x in y=2*(-5/3)+2 y=-10/3+2 y=-4/3

OpenStudy (anonymous):

okae ill check both of them thx guys